### v1

# Machine Learning Should Maximize Welfare, Not (Only) Accuracy

Abstract
Decades of research in machine learning have given us powerful tools for making accurate predictions. But when used in social settings and on human inputs, better accuracy does not immediately translate to better social outcomes. This may not be surprising given that conventional learning frameworks are not designed to express societal preferences—let alone promote them. This position paper argues that machine learning is currently missing, and can gain much from incorporating, a proper notion of social welfare. The field of welfare economics asks: how should we allocate limited resources to self-interested agents in a way that maximizes social benefit? We argue that this perspective applies to many modern applications of machine learning in social contexts, and advocate for its adoption. Rather than disposing of prediction, we aim to leverage this forte of machine learning for promoting social welfare. We demonstrate this idea by proposing a conceptual framework that gradually transitions from accuracy maximization (with awareness to welfare) to welfare maximization (via accurate prediction). We detail applications and use-cases for which our framework can be effective, identify technical challenges and practical opportunities, and highlight future avenues worth pursuing.
1 Introduction
As the influence of machine learning on our lives continues to grow, there is high hope that this will prove to be for the better. Given the unmatched ability of machine learning to make accurate predictions, we have, in principle, much reason for optimism. Accurate predictions have potential to improve the choices we make for ourselves, from the mundane (e.g., what to buy or where to eat) to highly consequential (e.g., what to study, how to invest, or even who to date). This can explain why machine learning has become the backbone of most recommendation systems, media services, online marketplaces, and social platforms. Better prediction can also provide better support for decisions made about us, such as which medical treatment to apply, when to offer financial aid, or who to hire. Learned models are therefore making their way, slowly but surely, into more conventional social domains such as health care, education, finance, transportation, law, and even government. Given how learning is increasingly shaping how we communicate, express ourselves, and are productive—its influence and integration are only likely to grow. This requires us to think carefully about the role we envision for machine learning in society, as it is forming now, and as we would like it to be.
If we accept that machine learning has the capacity to be socially beneficial, then ideally, we would like it to help in improving outcomes for everyone. But economic theory casts doubt as to the feasibility of such an aspiration. A basic economic truism is that whenever people benefit from something, that something will become a scarce resource, over which people will then contend. Our key point, which we will argue throughout, is that learning in social context is inherently, and inevitably, about how some form of limited resource is allocated; we will give various examples of such. When there is more demand than resources, the inevitable reality is that not everyone can get a piece of the cake. The important implication is that learning algorithms become the determinants of allocation, finding themselves in a position of making decisions about who gets what (and who doesn’t). For some tasks this is intentional, but more often it is not.
In some cases, predictions support downstream decisions in which resources are clearly scarce, such as in university admissions or job hiring, or in which risk must be hedged, such as loan approval or insurance. In other cases, scarcity can be more nuanced, but is nonetheless a substantial factor; for example, digital goods are technically of unlimited supply, but which recommended items are eventually purchased, and from which producers, is naturally restricted by the (bounded) choices of users. We posit that the issue of limited resources and their allocation holds much more broadly, and applies to most predictive tasks that takes place in social context—even when there are no explicit resources, or when the true objective is to simply maximize accuracy.
The problem is that conventional learning frameworks are designed to make predictions, not allocate resources. Thus, as we pursue ever-improving accuracy, we are possibly missing out (or even deteriorating) the ability to make effective and beneficial allocations. We believe that many instances in which the use of machine learning has led to undesired outcomes, even if unintentional, can be explained by this shortcoming. As we argue, the main issue with limited resources is not their limitation per se, but the fact that people have interest in securing those resources for themselves. This creates competition either between the system and its users (if they have differing interest), amongst users themselves (directly or indirectly), or both. The crux of conventional learning tools is that they fail to acknowledge the role of human agency and its implications on learning outcomes.
Our goal is to promote the study of learning algorithms that are aware of the economic context in which they operate, and are explicitly designed to promote favorable social outcomes—by accounting for resource scarcity and human agency, and how they relate. Towards this, our main thesis is that learning frameworks should be designed to innately support a notion of social welfare. In economics, welfare quantifies the overall benefit to agents in an economic system, and questions regarding the estimation, maximization, and distribution of welfare are central to the analysis of any such system. Drawing on ideas from the field of welfare economics, we propose to adopt its main principles—namely efficiency and equity—and adapt them to the purposes and needs of current machine learning practice. We discuss the potential merits of this approach, the technical and conceptual challenges it presents, the necessary steps to ensure its successful adoption, and the limitations that will likely still remain.
1.1 Setting and scope
For concreteness, we focus on the fundamental learning task of maximizing predictive accuracy by training a model on a sample set of labeled data. However, all the ideas described here are easily applicable to other learning paradigms, including today’s generative models. We follow the conventional supervised learning paradigm and assume data is in the form of feature-label pairs $(x,y)$ sampled iid from some unknown distribution ${D}$. The data we consider represents humans; we will typically think of inputs $x$ as describing individuals (e.g., loan application, resume, movie viewing history, past consumer choices), labels $y$ as describing outcomes of interest (e.g., loan return or default, job qualification, movie rating, satisfaction), and ${D}$ as representing the population.
Although our goal is to discuss how learning can promote social good, we will insist that accuracy remains an integral part of the learning objective. This is both because accuracy has value in and of itself, and because predictive methods are, and likely will continue to be, a primary tool used in practice. Thus, we seek to retool the conventional supervised learning framework towards the goal of welfare maximization through the use of accurate predictions. This will allow us to capitalize and build on existing tools, knowledge, and practice. Because we consider limited resources and their (sometimes implicit) allocation, our discussion must extend to consider policies operating on the basis of predictions, in the style of prediction policies of Kleinberg et al. (2015).
1.2 Limitations of current practice
It is tempting to hope that more data, more compute, and more sophisticated learning algorithms will pave the way to a better tomorrow. But merely improving predictions, whether to match or go beyond human capabilities, still presents inherent limitations. Consider an example:
A school receives funding for providing students with targeted aid intended to help them pass an exit exam that is required for graduation. To decide who will be selected for the program, the school obtains past data on students that participated in earlier programs and trains a model to predict the likelihood of the aid being helpful. It then uses the model to determine which new students to admit.
On its face, this approach seems reasonable—but of course things are not so simple: the school may wish to maximize the impact of aid, but doing so by maximizing accuracy falls short on some important aspects. Consider the following: [^1] Other issues are apparent in the example, e.g., distribution shift or causal considerations—but these are orthogonal to our concerns.
  • Although the learning objective gives equal importance to all students, this does not imply that all predictions will be equally accurate. Since borderline cases are tougher to decide, decisions could improve by giving them precedence.
  • Students that are selected may not necessarily choose or agree to participate. The program might find itself with very few students and unused resources.
  • Contrarily, students that are interested in receiving aid are likely to compete for acceptance, given that space is limited. This can bias selection towards students that were able to ‘appear’ more qualified (or that cheated).
  • The above approach selects students on the basis of whether aid will be effective. It does not ask how beneficial it is for each of the students to pass the exam.
  • Maximizing accuracy is used here as a proxy for maximizing the number of successes. This makes no consideration of who receives aid, nor of how overall benefits from aid are distributed across the body of students.
The example shows that while accuracy goes some way towards welfare, there is still an inherent tension between them. Luckily, this tension has structure, which reveals several ‘blind spots’ of current machine learning practice. These include: (i) limited resources, (ii) system decisions, (ii) human agency, and (iv) welfare considerations. Our framework intends to capture this structure and exploit it for promoting welfare through predictive machine learning.
1.3 What has been done, and what is still missing
The issue of aligning learning with societal preferences is of course not new, nor is the introduction of social welfare (Shirali et al., 2024; Perdomo, 2024). We believe however that the current literature requires a unifying formal framework for this notion. Several subfields of machine learning have made advances on some fronts, but remain to lack in others. A prime example is the literature on fairness in machine learning (Dwork et al., 2012), which clearly aims to advance some form of social good, and for this has received much attention. One concern that has been voiced is that fairness focuses on making things equal, but is unable to express the benefit that results from such efforts (Heidari et al., 2018; Hu and Chen, 2020). From our perspective, fairness takes one step towards acknowledging that something is of limited supply (e.g., positive predictions), but this is not made explicit. More importantly, fairness constraints do not account for agency: they do not ask what users want, nor how much they stand to lose or gain. Another issue is that such constraints do not consider user actions; for example, predictions for school admissions may satisfy demographic parity, but if members of the minority group do not choose to apply initially, then this guarantee is broken (Horowitz et al., 2024).
One line of research that is close to ours is that of strategic learning. This includes topics like strategic classification (Brückner et al., 2012; Hardt et al., 2016) and performative prediction (Perdomo et al., 2020) which have drawn much recent interest. Strategic learning acknowledges and explicitly models user agency, and in some cases, scarce resources. And while some works discuss social aspects of learning (e.g., Milli et al., 2019), the focus remains primarily on maximizing accuracy (and to some degree, notions of stability at equilibrium). The idea of welfare is not a driving consideration, nor is it an integral part of these frameworks.
We believe there is need for an umbrella framework that ties these loose ends and incorporates them as special cases. Such a framework will also be useful for pointing out what is missing in current solutions, and what broader questions still needs addressing. Welfare in economics is rarely something that is forced on a system; rather, it should emerge as the solution to the question—‘how can we appropriately align incentives’? Another goal is therefore to provide social planners and policymakers a handle for steering learning toward preferable social outcomes, as they deem appropriate.
1.4 Paving the path to welfare-aware machine learning
Machine learning offers powerful tools for training accurate predictive models, but the above limitations suggest that this might not suffice if we wish to solve more stressing societal problems. Not by chance, social welfare offers a complementary perspective that enables such considerations. We believe that current machine learning is missing, and could much benefit from, incorporating a notion of welfare.
Position: In social contexts, machine learning should enable us to improve our collective well-being. This requires a learning framework that is capable of expressing, and provides means to optimize, the welfare of its users. To promote this goal, we should exploit the power of learning to obtain accurate predictions—not shy away from it. But accuracy cannot be the only consideration; to support constructive social outcomes, learning must: (i) explicitly consider resources, scarcity, and allocations; (ii) account for human agency by modeling what users want, know, and do; and (iii) adhere to well-specified social welfare objectives designated by a social planner.
Throughout the paper we will make these ideas clearer, more precise, better grounded, and placed in context.
Paper outline.
Our main thesis is that machine learning can benefit from the ideas and tools that welfare economics has to offer; Sec. 2 discusses its main principles, and highlights the potential for synergy with machine learning. In Sec. 3 we present our proposed framework for welfare-maximizing machine learning and details its three orders. Sec. 4 presents alternative views, and Sec. 6 gives concluding remarks and a look to the future.
2 Welfare economics and what it can teach us
Welfare economics is the subfield of economics that is concerned with the characterization, evaluation, and maximization of social welfare in economies and societies. The main principles of welfare economics date back to Adam Smith (1759, 1776), but its formal foundations were laid out only a century later by notable neoclassic economists such as Edgeworth (1881), Marshall (1890), Pareto (1906), and Pigou (1920). Modern welfare economists include influential figures such as Hicks (1939), Arrow (1951), Sen (1970), and Stiglitz (2012)—all of which received the Nobel Prize for their contributions to this field. Given its rich history, we believe that machine learning has much to gain from adopting ideas and perspectives from this well-established discipline.
2.1 Welfare economics: crash course
The distribution of wealth.
The main question welfare economics asks is: how should wealth be distributed across individuals in the economy? Under the working hypothesis that resources (and therefore wealth) are limited, the main object of interest in welfare economics is the Pareto front—the set of all possible economic states in which no individual can be made better off without making things worse for another (see, e.g., Johansson, 1991). In terms of welfare, there are two main considerations for policymakers:
  1. 1.
    Efficiency: How do we reach a Pareto state?
  2. 2.
    Equity: Of all Pareto states, which are preferable?
Efficiency requires an ability to optimize economic outcomes to a point where social benefit is maximal (in the Pareto sense). Markets are a classic example of how the actions of many self-interested agents can combine to produce efficient outcomes (Arrow and Debreu, 1954). However, there are typically many states that are maximally beneficial, but that differ in how benefits are distributed across individuals; in regard to this, markets are mostly silent. As such, equity makes a statement about the relative preference ordering over all possible states, and considers means for steering towards preferable ones. A canonical example is income distribution: all governments likely seek higher overall income (efficiency), but may disagree about whether high inequality should be permitted or suppressed (equity).
Social welfare functions.
In welfare economics, the primary tool for defining and promoting equity is the social welfare function, which ranks or scores all possible economic states. Our focus will be on the common choice of cardinal (i.e., real-valued) welfare functions that take the form of an expectation over weighted individual utilities: [^2] The literature on social welfare functions is of course rich and diverse. Here we present a basic and simplified formulation which suits our purposes. Curious readers are referred to (Adler, 2019) for a more thorough discussion of possible alternatives.
${\mathtt{welfare}}({\pi};w)=\mathbbm{E}_{(x,y)\sim{D}}{\left[{w(x)u(x,y;{\pi})}\right]}$ (1)
where ${\pi}$ is a policy (e.g., rules deciding which resources are allocated to whom) and $u(x,y;{\pi})$ is the utility of user $x$ under policy $\pi$, given $y$. Notably, $w(x)$ is a weight function that the planner chooses: this defines the desired direction for overall improvement (i.e., efficiency) by balancing the importances of individual outcomes (i.e., equity). [^3] More generally, weights $w(x)$ can even depend on the utility $u$ of $x$. This allows to express welfare using broader functional forms (e.g., $\min$ instead of sum) or relative measures of inequality (e.g., Gini or Theil). For our purposes, the simpler form $w(x)$ suffices. When the policy ${\pi}={\pi}_{h}$ is guided by a predictive model $h$, we will write ${\mathtt{welfare}}(h)$ to mean ${\mathtt{welfare}}({\pi}_{h})$.
The social planner.
Equity is inherently a subjective notion that requires making value judgements. Social welfare functions make it possible to formally express these by setting appropriate weights. In welfare economics, weights are designated by a social planner—either a real or fictitious entity that represents societal preferences. A social planner can set weights to aid low income individuals; implement affirmative action towards some social group; or ensure that all individuals obtain some minimal level of utility. Concrete examples include using weights to prioritize certain subgroups (Björkegren et al., 2022) or balance different objectives (Rolf et al., 2020). The simplest weighing scheme is of course using uniform weights, i.e., $w(x)=1$ for any $x$. But note even this makes a statement, which is that individuals should be weighted by their utility; this is known as ‘utilitarian welfare’. Hence, from the perspective of welfare economics, any objective that optimizes a (non-weighted) average—such as accuracy in machine learning—is in effect making a statement about how value should be distributed.
Human agency.
Welfare economics makes explicit the idea that individuals have agency. Intuitively, this states that individuals (i) want things, (ii) know things, and (iii) act – to get what they want, using what they know. These notions are formally accounted for by modeling utility functions (want), private information (know), and decision-making, e.g., rational or behavioral (act). Any policy that aims to advance welfare must take these into account. Often this requires the planner to make additional efforts, such as to elicit preferences, create incentives for truthful reporting, or infer how users will respond to different policy choices. These are challenging, but give the planner power: if incentives can be aligned, then it becomes possible to harness the willingness of users to invest effort for improving outcomes for all; consider public goods and services, crowdfunding platforms, open-source software, and collaborative knowledge bases.
2.2 Connections to machine learning.
We believe that machine learning has much to gain from adopting a welfare perspective: when inputs represent humans, it becomes possible to promote overall social benefit (efficiency), and imperative to consider its distribution across the population (equity). Current tools already push forward on these ideas, but only to a limited extent. One reason is that standard learning objectives are notoriously underspecified; this has implications on e.g. robustness (D’Amour et al., 2022), explainability (see Rudin, 2019), and fairness (e.g., Rodolfa et al., 2020; Coston et al., 2021; Black et al., 2022). But underspecification also presents an opportunity for steering outcomes toward socially beneficial states. Consider how the idea of model multiplicity (Breiman, 2001; Marx et al., 2020; Hsu and Calmon, 2022), i.e., that there is typically a large set of (approximately) optimal models, connects to the notion of an efficient Pareto front. The challenge lies in how to provide a social planner effective means to choose a preferred operating point. In some cases, this will be diffucult; in others, it may be achievable with tools as simple as regularization—if chosen appropriately (Levanon and Rosenfeld, 2021). In terms of limited resources, machine learning already offers many relevant tools, such as constraints on cardinality (e.g., top-$k$ prediction) or error rates (e.g., precision and recall). Also relevant are tools from cost-sensitive (e.g., Elkan, 2001) and decision-focused learning (Mandi et al., 2024). But to be effective, these must be adapted to account for agency: how users report information, how they act in response to the learned classifier, and how they interact with each other. This idea will be key to our framework, as discussed next.
3 Three Orders of Welfare-Maximizing ML
Figure 1: Proposed framework: The three orders of welfare-maximizing machine learning.
Our framework presents a hierarchy of problem formulations for welfare-maximizing learning, organized into three ‘orders’ of gradually increasing complexity. The orders are built in a bottom-up fashion: The lowest order (Order 0) coincides with the standard objective of maximizing predictive accuracy, but provides a novel welfare perspective suited for social tasks. Each higher order then builds on and generalizes the one below it, this by adding to the objective another layer of economic complexity: first focusing on system decisions (Order 1), and then on user choices (Order 2). As such, our framework is based on accuracy maximization at its core, but enables—and in fact requires—to explicitly model resources, allocations, and agency. It also requires to specify the role predictions play in shaping social outcomes. An illustration of the framework and its orders is given in Fig. 1.
We begin with a welfare interpretation of supervised learning, and then proceed to discuss the three framework orders. Further prospects and challenges are discussed in Appx. 5.
Supervised learning from a welfare perspective.
Given labeled training data of pairs $(x,y)$ sampled iid from an unknown distribution ${D}$, the goal in supervised learning is to find a function $h$ from a chosen class $H$ whose predictions $\hat{y}=h(x)$ are accurate on future examples in expectation:
$\operatorname*{argmax}\nolimits_{h\in H}\mathbbm{E}_{{D}}{\left[{\mathds{1}{\{{y=h(x)}\}}}\right]}$ (2)
Since in our setting examples $(x,y)$ are associated with individuals in the population ${D}$, then from a welfare perspective, we can think of Eq. (2) as prescribing a particular social welfare function—a special case of Eq. (1)—in which: (i) utility to users derives from accurate predictions, (ii) all users share the same utility function, $u(x,y;{\pi}_{h})=\mathds{1}{\{{y=h(x)}\}}$, (iii) possible outcomes include correct and incorrect predictions, hence (iv) the policy is degenerate, $u(x,y;{\pi}_{h})=u(x,y;h)$), and finally (v) weights are uniform, $w(x)\equiv 1$.
From accuracy to welfare maximization.
Despite an apparent ‘neutrality’, the above reveals that accuracy maximization in fact makes an (implicit) statement about social preferences. This entails a particular notion of equity—one that derives operationally from the predictive task at hand, rather than from a planned or designated social goal. But learning is of course not bound to maximizing objectives only of the form of Eq. (2). If our true goal is to maximize welfare, then we should modify the objective to support this goal. For example, if users differ in how much they benefit from accurate predictions, then we can plug in an appropriate utility function $u$; or if there are exogenous constraints on predictions, then we can express these via the policy ${\pi}_{h}$. Such steps are certainly possible, but introduce challenges beyond those found in standard machine learning tasks. Towards this, we pave a path that gradually transforms Eq. (2) to support increasingly complex economic considerations.
3.1 Order 0: Allocating Accuracy
Accuracy will be beneficial to users in any a platform or service that relies on predictions; consider recommendation systems, online market platforms, financial aid tools, or diagnistic health services. We argue that in such cases, and as long as learning is prone to some level of error (which is plausible), then accuracy itself is a limited resource—simply because not all users can obtain the same level of accuracy. Thus, the choice of classifier becomes, in effect, a decision on which users will be ‘allocated’ more accurate predictions, and which less. This then begs the question: who should be allocated an accurate prediction? One answer is users who most need it, or who benefit mostly. In the reasonable case where there is natural variation in individual utilities, we can to optimize a utility-weighted accuracy objective:
$\operatorname*{argmax}\nolimits_{h\in H}\mathbbm{E}_{{D}}{\left[{u(x,y;\hat{y})\cdotp\mathds{1}{\{{y=h(x)}\}}}\right]}$ (3)
where $u(x,y;\hat{y})$ is the utility for user $x$ given prediction $\hat{y}$. This transforms Eq. (2) into an explicit utilitarian welfare objective in which individualized utilities derive from prediction correctness: $u(x;{\hat{y}})$ if ${\hat{y}}=y$, and zero if not. Conditioning $u$ also on predictions ${\hat{y}}$ enables to express utilities that depend on predictions themselves (beyond their correctness); for example, if users benefit more from positive predictions, as in loans, hiring, admissions, etc.
Agency.
While apparently simple, optimizing Eq. (3) poses significant challenges once we accept that users have agency. If users seek to promote their own interests, and have control of information that is important to the system, then their choices regarding how and what to report can significantly affect learning outcomes. For example, consider that if we simply ask people to report their utilities, then each individual will be incentivized to report the highest possible value, making all reports completely uninformative. Another example is that users may be prone to manipulate their features if this helps them secure favorable predictions, as is the case of strategic classification (Hardt et al., 2016).
Challenge: Optimizing welfare objectives.
Maximizing welfare through learning presents new algorithmic challenges. One task is to identify learners that optimize a given social welfare function—intrinsically, a question of identifying a desirable point on the Pareto front. Another task is to generate all points on the frontier (see, e.g., Navon et al., 2021). One possible approach is to cast Eq. (3) as a problem of distribution shift. Here shift can stem from several factors, including the welfare function weights, user utilities, and user inputs. But agency means that utilities can be misreported and features manipulated; as a result, the way in which the distribution shifts becomes dependent on the choice of classifier, indirectly through how users respond. This makes Eq. (3) a challenging instance of model-dependent distribution shift (Drusvyatskiy and Xiao, 2023). Here the interaction between the system and its users can be modeled as a Stackelberg game (Chen et al., 2020). This perspective has been useful for questions on learnability, e.g. via generalization bounds that rely on strategic VC analysis (Zhang and Conitzer, 2021; Sundaram et al., 2023) or that accommodate feature-dependent utilities (Pardeshi et al., 2024). Nonetheless, many important open questions remain.
3.2 Order 1: Incorporating System Decisions
Generally, predictions are useful for the system if they aid in making better decisions about uncertain outcomes. Since decisions in social settings often directly prescribe how to allocate limited resources to users, a useful next step is to encode them explicitly into the objective. We formalize this idea by modeling a system that makes decisions about users (e.g., who to hire) through predictions (e.g., resume screening). Consider a user $x$ with label $y$, and denote by $a$ the action the system takes (e.g., hire or not). Let ${r}(a,y)$ be the reward for the system on this user given action $a$ (e.g., the quality of the candidate, if hired). The direct dependence of ${r}$ on $y$ means that if we know $y$, then we can write the optimal action as $a^{*}={\pi}(y)$ for some policy ${\pi}$. Since $y$ is generally unknown, the common approach is to replace $y$ in ${\pi}$ with a prediction ${\hat{y}}=h(x)$, denoted $a={\pi}(h(x))={\pi}_{h}(x)$, in hopes that better predictions translate to better decisions. Such policies, referred to as prediction policies (Kleinberg et al., 2015), are appropriate when the uncertainty in $y$ is a stronger factor for outcomes than potential causal effects (we discuss this distinction further in Appendix 5.4). In this setting, choosing to work with prediction policies gives reason for the system to optimize $h$ for accuracy (Eq. (2)).
From predictions to actions.
We consider cases where there is a global restriction on the set of all actions $a$. These can express, for example, a limited number of available jobs (via cardinality constraints), a total sum of funds that a bank can lend (knapsack constraints), regulation on the amount of financial risk an insurer can take (bounded expected risk), or a limit on the number of posts a social platform can block to still enable free speech (lower-bounded rates). Given a set $A$ of feasible actions, the objective can be written as:
$\operatorname*{argmax}\nolimits_{h\in H}\mathbbm{E}_{{D}}{\left[{{r}({\pi}_{h}(x),y)}\right]}\quad\text{s.t.}\quad{\pi}_{h}({D})\in A$ (4)
where ${\pi}_{h}({D})$ is the set of actions over the entire population.
Agency.
Machine learning has many tools for coping with constraints. For example, if there is a quota on the number of available interview slots, then top-$k$ classification (e.g. Lapin et al., 2015; Petersen et al., 2022) or ranking (e.g. Cao et al., 2007) seem like adequate solutions. The crux, however, is that such methods do not account for agency: since users seek to be ‘allocated’ (e.g., get the interviews), they will likely try to present themselves as relevant (e.g. by tweaking their resume). A key point is that since users contend over resources, this introduces dependencies into their actions and subsequent outcomes. Such behaviors are ubiquitous in economics (e.g., costly signaling, adverse selection), but have been underexplored in machine learning.
Social welfare.
In terms of welfare outcomes, we make a distinction between two settings of interest:
  • An aligned setting in which ${r}={\mathtt{welfare}}$, and so the interests of the system align with societal preferences, e.g., as for government and nonprofit organizations. Note Eq. (3) becomes a special case of Eq. (4) when utility derives from accuracy and $A$ does not impose restrictions.
  • A misaligned setting in which ${r}$ can be at odds with welfare, for example if it concerns revenue or user engagement, as is more common in commercial settings. Here a social planner is needed to incentivize the learner to account for welfare outcomes, as we detail in Sec. 5.1.
Challenge #1: Decisions and externalities.
At Order 0, user responses are generally made independently, and so the learning objective decomposes over examples. [^4] Note this is a key assumption in standard strategic classification. But once the system makes decisions under constrained resources, user behavior can become intricately dependent. Consider for example admissions under a limited quota: here applicants cannot simply pass a fixed acceptance threshold, but instead, must surpass other applicants—i.e., the bar adjusts according to competition. The tension for both decision-makers and applicants tends to focus on ‘marginal’ students; complex gaming behavior among such candidates is well-documented in the education literature (Bound et al., 2009). In economics, inter-user dependencies are known as externalities. One way to model how learning creates externalities and coordinates user behavior is by casting the problem as a Stackelbeg-Nash game (Liu, 1998), where the system plays first, and users respond collectively by playing an induced simultaneous game (with externalities). Not much work has studied learning in this challenging setting.
Challenge #2: Misaligned system and user objectives.
As we note, generally we cannot expect the reward for the system to align with user interests. If no external forces intervene, then the challenge is to learn a decision policy that is effective at the Stackelberg-Nash equilibrium. This necessitates learning objectives that can express and anticipate how users will collectively respond to a deployed model. Contrarily, if there is regulation (e.g., on welfare or risk), then the system’s decision space becomes much more limited. An interesting question then is how to align incentives through regulation. For example, would requiring the system to optimize a weighted sum of accuracy and welfare help to promote equity? And if so, which weights, and by what means?
3.3 Order 2: Enabling User Choices
Even if a system has the capacity to make decisions about users, often those very users will also have some say regarding final outcomes. For example, even in hiring and admissions, candidates much first choose to apply, and if selected, choose to accept. Thus, both reward for the system and utility to users depend on outcomes that result from the interplay of system decisions and user choices. Similarly to system decisions, we will model user choices as also depending on predictions, or more generally on the learned model $h$. Let ${c}_{{\pi}_{h}}(x)$ be the choice of user $x$ in response to $h$. Incorporating user choices into the system’s objective gives:
$\operatorname*{argmax}_{h\in H}\mathbbm{E}_{{D}}{\left[{{r}({\pi}_{h}(x),{c}_{{\pi}_{h}}(x),y)}\right]}\quad\text{s.t.}\quad{\pi}_{h}({D})\in A$ (5)
where ${r}$ now depends jointly on system decisions ${\pi}_{h}$ and user choices ${c}_{{\pi}_{h}}$ through the learned predictor $h$.
Agency and resources.
When users make choices, this makes them a part of the allocation process—but also the reason for scarcity. Here we outline three scenarios of interest:
  • Self-selection: For most systems that makes decisions about users, those users must first choose to join the system; consider hiring, admissions, medical programs, educational aid, and welfare benefits. In econometrics, the choice of users to participate (or not) based on the utility they expect to obtain is known as self-selection (Roy, 1951). Since self-selection in aggregate determines the user population, once user decisions depend on the learned model, learning obtains the power to shape the population’s composition—a power which should not be taken lightly. Although self-selection has a long history in economics, it is only now beginning to draw attention in machine learning research (Zhang et al., 2021; Cherapanamjeri et al., 2023; Horowitz et al., 2024).
  • Matching: Prime examples in which user choices drive economic outcomes are recommendation systems and online market platforms. These typically include a single learner whose role is to match users with relevant items (Jagadeesan et al., 2023b); once matched, it is the choices of users that ultimately determine allocations. Since matchings are often based on predictions, their quality determines outcomes for users, suppliers, and the system. This makes them unique in relation to conventional two-sided markets (Rochet and Tirole, 2003). In modern human-facing systems, key resources primarily include the attention, time, and budgets of users. This calls for careful and responsible modeling of user choice behavior (Kleinberg et al., 2024).
  • Competition: For many tasks and services of interest, users can now choose between multiple platforms or providers. This creates competition: once providers aim to maximize their market share, users themselves become the scarce resource, over which providers compete (Ben-Porat and Tennenholtz, 2019; Jagadeesan et al., 2023a; Yao et al., 2023, 2024). When the benefit of users depends on the quality of predictions, competition will revolve around which provider learns more accurate models—and for which users. This creates a novel ‘market for accuracy’, in which learning can have much impact on efficiency and equity.
Social welfare.
A canonical property of classic markets is that they coordinate the behavior of many self-interested agents in a way which can naturaly lead to welfare maximization (e.g., Arrow and Debreu, 1954; Shapley and Shubik, 1971). A key question in our context is whether this emerges also in markets where coordination is mediated by predictions (Nahum et al., 2024). Given the growing concerns regarding how recommendation systems may drive polarization, echo chambers, informational barriers, economic inequity, and unhealthy usage patterns—the answer is likely negative. Welfare economics may be helpful in posing the question of why this happens as one of market failure. This provides tools for uncovering the mechanisms underlying failure, such as negative externalities, public goods, information asymmetry, market power and control (e.g., by monopolies), or collusion, as they manifest through learning.
Challenge #1: Escaping echo chambers.
Since its early days, the task of recommending content has been treated as a pure prediction problem. But accumulating evidence of its likely ill effects has motivated a search for more viable alternatives. A major issue is the reinforcing nature of accuracy-driven recommendation: users that are recommended certain items become ‘locked in’ on similar content through a positive feedback loop of indefinite model retraining, choice behavior, and recommendation. One aspect of this loop that is often overlooked, and relevant to our context, is that recommendations also incentivize the creation of new content by exposure-seeking creators. Recommendation is essentially a problem of matching demand (of users) and supply (of new content) over a bi-partite user-content graph; thus, if we wish to prevent the formation of echo chambers and filter bubbles, we should find ways to exploit this structure to promote socially-beneficial outcomes. Some examples include incentivizing exploration (Mansour et al., 2020) or the creation of diverse content (Eilat and Rosenfeld, 2023; Yao et al., 2024), but many other paths are possible.
Challenge #2: Learning to compete.
Once predictions are given as a service, it is only natural that multiple providers will compete over who can give users better predictions. One example are housing market platforms such as Zillow and Redfin which compete over who provides better pricing recommendations. To some extent, another example are LLMs who compete over better text generation. Competition is driven by the fact that users are free to choose a provider, but that choices are costly (e.g., time, premium features, opportunity costs). Thus, although users gain from accuracy, optimizing expected accuracy is likely not a good strategy for providers. As a result, welfare could suffer. Welfare in such markets will be high if the market is efficient. This requires a planner that can steer competing learners towards a favorable Nash-Stackelberg equilibrium, where users follow the simultaneous moves of multilpe systems.
4 Alternative Views
While it is easy to agree that learning systems should be designed to enable the promotion of social good, there will likely be disagreement as to how. We propose to adopt the perspective of welfare economics, but there certainly exist alternative viewpoints and complementary approaches.
Give accuracy time.
One perspective is that if we give data enough time to accumulate and new methods enough time to improve, then machine learning will organically overcome the challenges we discussed. One example to draw on is how despite many advances in optimization, the simple gradient descent algorithm still drives most modern tools. Another is how large language models have demonstrated that simply predicting the next word with sufficient accuracy and on enough data gives rise to emergent phenomena far beyond this basic task. Our position is that limited resources is an inherent problem of any social system, whether technology-driven or not. We believe that scarcity should be addressed explicitly—but of course we may be proven wrong.
Divide and conquer.
Even if machine learning as a standalone solution does not suffice, one could argue that an economic approach can be applied on top of existing learning tools, rather than integrated within them as we propose. Hence, learning and policy can be advanced independently and combined only later. This is reasonable, and independent efforts and application will likely be required regardless of whether direct integration works or not. But there is increasing evidence that this will not suffice; in fact, the field of fair machine learning rose in response to the clear need for embedding social considerations within the learning objective itself. Advances in the study of fairness in learning have also shown that fairness constraints alone cannot guarantee equity, such as when learning effects accumulate over time (e.g., Liu et al., 2018) or as a result of strategic user behavior (Horowitz et al., 2024). We take these to suggest that the novelty in the interface between learning and economics requires a wholistic approach specialized for this intersection.
Welfare without welfare.
Welfare economics is not the only approach for reasoning about and facilitating welfare, nor is it free from issues and limitations in itself. Criticism includes its subjective nature; the need to measure and compare utility across individuals; the emphasis on cardinal rather than ordinal utilities; the reliance on assumptions of rational behavior; the susceptibility to externalities and other sources of market failure; the need for a centralized social planner entity; and challenges in policy evaluation. Other schools of thought in economics offer alternatives: For example, Sen’s capabilities approach (Sen, 1999) focuses on ensuring people are capable of achieving what they seek, rather than the value of what they obtain. Within machine learning, there been have calls for alternative approaches as well, such as to ‘democratize’ the issue of alignment using social choice theory (Conitzer et al., 2024; Ge et al., 2024; Fish et al., 2024). We view these as complementary to ours, and believe there is merit advancing welfare in machine learning simultaneously along several fronts.
5 Further Prospects and Challenges
Sec. 3 lays out the main goals and principles of our framework, along with possible avenues to pursue within each of its three orders. But there are further opportunities, as well as challenges, that lie beyond the framework’s core and are of both interest and significance. In this section we highlight some of these directions.
5.1 Regulation: Social planner, revisited
Our perspective of a social planner so far has been that of a useful construct for defining the social preferences that should guide learning. As such, we have mostly discussed the question of how to learn in a way that promotes or aligns with given social preferences. In a sense, this depicts the social planner as acting before learning. But there are additional roles a social planner can play during or after learning, actively and through various mechanisms. For example:
  • Monitoring. One defining characteristic of data-driven systems is that they hold a distinct informational advantage over their users. In economics, information asymmetry is a well-known source of hazard for markets, with two prominent issues known as moral hazard (Holmström, 1979) (arising from asymmetric information on decisions) and adverse selection (Akerlof, 1978) (arising from asymmetric information on system states). Monitoring can help to mitigate this by identifying what information gives the system an unfair advantage, and then channeling this information to users. This provides an operational motivation for transparency, with the definitive purpose of improving welfare.
  • Subsidizing. When resources are limited, their allocation often depends on the wealth or assets of individuals. This presents an opportunity to intervene by injecting funds into the system in a way that balances outcomes, e.g., via targeted subsidies. Subsidies can either be given in response to a learned classifier, or determined by policy prior to learning. Since outcomes in our case depend on predictions, this presents a unique opportunity for a social planner to intervene by directly providing users with useful ‘features’, thus bypassing the need for a monetary mechanism. An interesting question is then how to design learning objectives that operate under such subsidies.
Other roles include: setting standards, licensing, auditing, mandating disclosure and transparency, and stress-testing.
5.2 Inverse tasks: policy preference elicitation
A key point we have made throughout is that any prediction policy advances some social preferences. Until now, our focus was on learning to maximize welfare under a specified social welfare function. But we can also ask the inverse question: given a policy—what social welfare function is it optimizing? This can be useful for revealing the preferences of general, unspecified policies, and for evaluating whether a policy that has been specified truly achieves its goals. One way to approach this would be to use tools from econometrics, e.g. as in (Björkegren et al., 2022; Pardeshi et al., 2024). But is also interesting to ask whether ideas from inverse problems in machine learning can be adopted.
5.3 Evaluation
A main driving force of modern machine learning is the ability to evaluate the performance of different methods effectively and consistently under a wide range of settings and conditions. Unfortunately, such privileges are not possible when learning in social settings. The challenge of evaluation is shared by all fields at the intersection of learning, economics, and human behavior, but welfare-maximizing machine learning presents its own unique difficulties. For example, even the basic task of measuring or estimating social welfare is far from trivial. One immediate challenges is the difficulty (and restrictions) of obtaining and working with representative and informative user data. Another is that, as we have stressed, welfare depends on outcomes, which are the product of decisions due to policy, and so evaluation requires either control of the policy or a way to estimate counterfactuals. Other than a few noteworthy exceptions (e.g., Björkegren et al., 2020; Haupt et al., 2023; Mendler-Dünner et al., 2024), most current works settle for ‘semi-synthetic’ evaluation that uses real data and simulated user behavior. Simulation is often useful as an intermediary step to drive the field forward (e.g., as in reinforcement learning), but the ultimate goal should be to enable seamless, realistic, in-the-wild evaluation.
5.4 Accounting for causal effects.
As we state in Sec. 1.1, our focus has been on prediction policies (Kleinberg et al., 2015) that specify actions as $a={\pi}_{h}(x)={\pi}(h(x))$. The key assumption of prediction policies is that knowing the true $y$ suffices for making an optimal decision; hence the motivation for replacing $y$ with the prediction ${\hat{y}}=h(x)$. But in realistic settings, $y$ is rarely the only source of uncertainty. Most decision settings have a casual aspect, meaning that actions $a$ can affect outcomes $y$. Hence, there is no ‘true’ $y$ on which we can draw, and the optimal action $a$ (or ‘intervention’) must include an understanding of its causal effects on $y$.
When learning in social settings, the issue of causality is further complicated by the fact that users have agency, and therefore act. Because learning determines policy, and policies shape users’ choice behavior, learning comes to have a second-order causal effect on $y$. The task of causal inference therefore attains two goals: understand the impact of decisions on outcomes directly, and indirectly through user choices.
Econometrics offers many tools for contending with questions of causal estimation. The main challenge is to make inference of causal effects on the basis of observed data—which is the common setting considered in learning. In cases where we can experiment, randomized control trials (or A/B tests) offer an alternative. Some of these ideas have been used in the literature on strategic learning (Miller et al., 2020; Shavit et al., 2020; Horowitz and Rosenfeld, 2023), mostly in relation to the interaction between the system and its users. A prominent point that concerns welfare is that a causal understanding can be used to improve outcomes—if incentives are set correctly (Kleinberg and Raghavan, 2020). Despite the importance of causal effects, we believe there is merit in first exploring question of welfare under policy predictions, as means to set the stage for more advanced questions concerning causality.
6 Concluding Discussion
Economics studies the tensions that arise from having limited means but unlimited wants. Our main thesis is that while the appetite for what machine learning can give us only grows, we do not yet have the tools to integrate the limitations inherent to any social system in a way that supports informed policy choices. Welfare economics provides general guidelines for designing systems that enable a social planner to express and promote outcomes that align with societal preferences, as they relate to scarcity in resources. We have identified ways in which we believe this perspective applies to, and can be useful for, the design of learning systems that operate in social contexts and with human agency. Scarcity in learned systems is often far more illusive than in classic economic settings that consider tangible goods. But nonetheless, it exists, and its implications bear much significance for the value that we can hope to obtain from using machine learning as individuals. Our proposal to augment conventional learning frameworks with a notion of social welfare serves as a first step towards this goal. But the road ahead is long and challenging, and requires careful deliberation of the societal role we envision for machine learning.
Impact statement
This paper proposes that machine learning can and should be used to maximize social welfare. In principle, and by construction, the impact of our proposed framework on society aims to be positive. But our paper also points to the inherent difficulties of identifying, and making formal, what ‘good for society’ is. We lean on the field of welfare economics, which has for decades contended with this challenge, for ideas on how the learning community can begin to approach this daunting task. However, even if these ideas are conceptually appealing, the path to practical welfare improvement presents many challenges—some expected, others unforseen. For example, we may specify incorrect social welfare functions; or we may specify them correctly but be unable to optimize them appropriately; or we may be able to optimize but find that our assumptions are wrong, that theory differs from practice, or that there were other considerations and complexities that we did not take into account. For this we can look to other related fields—such as fairness, privacy, and alignment in machine learning—which have taken (and are still taking) similar journeys, and learn from both their success and mistakes.
Any discipline that seeks to affect policy should do so with much deliberation and care. Whereas welfare economics was designed with the explicit purpose of supporting (and influencing) policymakers, machine learning has found itself in a similar position, but likely without any planned intent. On the one hand, adjusting machine learning to support notions, such as social welfare, that it was not designed to support initially can prove challenging. However, and as we argue throughout, we believe that building on top of existing machinery is a more practical approach than to begin from scratch. The necessity of confronting with welfare consideration can also be an opportunity—as we can leverage these novel challenges to make machine learning practice more informed, transparent, responsible, and socially aware.