### v2

# but Not by (Only) Maximizing Accuracy

Abstract
Decades of research in machine learning have given us powerful tools for making accurate predictions. This has made such tools appealing for use in social settings and on human inputs. Yet despite a lack of justification for why the generic approach of accuracy maximization can or should improve our collective well-being—and mounting evidence of likely adverse outcomes—it remains the widespread default. This position paper asserts that for machine learning to become socially beneficial, it must be embedded within a broader economic framework that explicitly aims to maximize social welfare. The field of welfare economics asks: how should we allocate limited resources among self-interested agents to maximize overall benefits? We contend that this perspective applies to many contemporary applications of machine learning in social contexts, and advocate for its adoption. Rather than disposing of prediction, we propose to leverage this forte of machine learning towards welfare maximization. We demonstrate this idea by portraying a conceptual framework that gradually transitions from accuracy maximization (with awareness to welfare) to welfare maximization (via accurate prediction). We detail applications and use-cases for which this framework can be effective, identify technical challenges and practical opportunities, and highlight future avenues worth pursuing.
Keywords Welfare-maximizing machine learning  $\cdot$ ML in social contexts
1 Introduction
As the influence of machine learning on our lives continues to grow, there is high hope that this will prove to be for the better. Given the unmatched ability of machine learning to make accurate predictions, we have, in principle, much reason for optimism. Accurate predictions have potential to improve the choices we make for ourselves, from mundane (e.g., what to buy or where to eat) to highly consequential (e.g., what to study, how to invest, or even who to date). This can explain why machine learning has become the backbone of most recommendation systems, media services, online marketplaces, and social platforms. Better prediction can also provide better support for decisions made about us, such as which medical treatment to apply, when to offer financial aid, or who to hire. Learned models are therefore making their way, slowly but surely, into more conventional social domains such as health care, education, finance, transportation, law, and even government. Given how learning is increasingly shaping the ways we communicate, express ourselves, and are productive—its influence and integration are only likely to grow. This requires us to think carefully about the role we envision for machine learning in society, as it is forming now, and as we would like it to be.
If we accept that machine learning has the capacity to be socially beneficial, then ideally, we would like for it to help improve outcomes for everyone. But economic theory casts doubt as to the feasibility of such an aspiration. A basic economic truism is that whenever people benefit from something, that something will become a scarce resource, over which people will then contend. Our key point, which we will argue throughout, is that learning in social context is inherently, and inevitably, about how some form of limited resource is allocated; we will give various examples. And when demand exceeds available resources, the inevitable reality is that not everyone can get a piece of the cake.
As learning algorithms increasingly influence who gets what and who doesn’t—whether by design or by consequence—it becomes imperative to understand the implications of learning under scarcity. The challenge is that conventional learning frameworks are innately designed to make predictions, not allocate resources. Thus, as we pursue ever-improving accuracy, we are possibly missing out on (or even deteriorating) the ability to make socially effective and beneficial allocations. The main challenge of allocation is not the limitation of resources per se, but the fact that people have interest in securing those resources for themselves. This creates competition either between the system and its users (when their interest differ), amongst users themselves (directly or indirectly), or with other stakeholders. Thus, a second crux of conventional learning is that it fails to acknowledge the role of human agency and decision-making through its implications on learning outcomes. We believe such economic grounds underlie (to some degree) many instances of machine learning leading to undesired (if unintentional) outcomes, including: perpetuated biases (Zou and Schiebinger, 2018) and group exclusion (De-Arteaga et al., 2019); Goodhart’s law (Manheim and Garrabrant, 2018) and gaming behavior (Hardt et al., 2016); content polarization (Bakshy et al., 2015), extremification (Ribeiro et al., 2020), and echo chambers (Cinelli et al., 2021); popularity bias (Abdollahpouri et al., 2017), exposure bias (Salganik et al., 2006), algorithmic bias (Kleinberg et al., 2018), and the Matthew effect (Perc, 2014) as some examples.
Rather than hoping for social outcomes to miraculously improve as a result of accuracy maximization, this position paper advocates for the study, design, and deployment of learning algorithms that directly account for the economic context in which they operate. Economic awareness does not require reinventing the entire learning pipeline; rather, we assert this can, and should, be achieved by augmenting existing predictive methodologies to support key economic notions, foremostly resource scarcity and human agency. But “better social outcomes” is an illusive concept that often evades proper definition; or when it is defined, remains constrained to a particular narrow formalism. To maintain generality and provide flexibility, we propose to build on the notion of social welfare as a core guiding principle. In economics, social welfare quantifies the overall benefit to agents in an economic system, and questions regarding the estimation, maximization, and distribution of welfare are central to the analysis of any such system. Adopting a welfare economics perspective offers many benefits: an established conceptual framework for reasoning about social outcomes, formal structure that enables effective integration with machine learning, and a practical toolkit for guiding system designers and policymakers.
Although the concept of welfare has been considered to some extent in machine learning (e.g., Rambachan et al., 2020), it has so far been limited mostly to works at the intersection of computational game theory, and in a restricted manner, to fairness in machine learning. As more and more ML sub-communities are finding themselves in a position where their work begins to have significant social impact, this paper calls for a more comprehensive interface with welfare: in theory and in practice, throughout the entire learning pipeline, and by any and all in the general learning community. Towards this, we suggest a conceptual framework consisting of three "orders" of problem formulations, targeting increasingly complex economic constructs and providing different levels of abstraction. We intend for our framework to serve as a roadmap for researchers interested in integrating welfare considerations into learning tasks of all types; to enable a bidirectional flow of ideas and methodologies between learning and economics; and to permit effective communication across the different learning communities. We discuss the potential merits of this framework, the technical and conceptual challenges it presents, the necessary steps to ensure its successful adoption, and the limitations that will likely still remain.
1.1 On the limitations of accuracy: an illustrative example
It is tempting to hope that more data, more compute, and more sophisticated learning algorithms will pave the way to a better tomorrow. But merely improving predictions, whether to match or go beyond human capabilities, still presents inherent limitations. Consider an example:
A school receives funding for providing students with targeted aid intended to help them pass an exit exam that is required for graduation. To decide who will be selected for the program, the school obtains past data on students that participated in earlier programs and trains a classifier to predict the likelihood of aid being helpful. It then uses the model to determine which new students to admit.
On its face, this approach seems reasonable—but of course things are not so simple: the school may wish to maximize the impact of aid, but doing so by maximizing accuracy as a proxy falls short on several important aspects. Consider the following: [^1] The example exhibits other issues, e.g., data shift and causality—which are important, but orthogonal to our goal.
  • Although the learning objective gives equal importance to all students, this does not imply that all predictions will be equally accurate. Cases may differ in their influence on final decisions, their difficulty of resolution, their impact on individuals, or their implications for the entire program.
  • Even if a student is selected, he or she may not necessarily choose or agree to participate. Without factoring participation, the program might find itself left with fewer or less qualified students.
  • Contrarily, students interested in receiving aid are likely to compete for acceptance, given that space is limited. This can bias selection towards students trying to ‘appear’ more qualified (or cheating).
  • Maximizing accuracy is used here as a proxy for maximizing the number of successes. This makes no consideration of who receives aid, nor of the distribution of benefits from aid across all students.
  • The above approach selects students on the basis of whether aid will be effective. It does not ask how beneficial it is for each of the students themselves to be admitted or to pass the final exam.
The example shows that while accuracy goes some way towards welfare, there is still an inherent tension between them. Luckily, this tension has structure, which reveals several ‘blind spots’ of current machine learning practice. These include: (i) limited resources, (ii) human agency, and (iii) policy considerations. Not by chance, these are precisely the building blocks of welfare economics (see Sec. 2).
1.2 Machine learning for social welfare
Machine learning offers powerful tools for training accurate predictive models, but the above limitations suggest that this might not suffice once learning encounters human reality. Social welfare offers a complementary perspective that enables such considerations. We believe that current machine learning is missing, and could much benefit from, incorporating a notion of welfare.
Position: Machine learning has the potential to improve our collective well-being, and should be used towards this purpose. But we cannot assume social good will spontaneously materialize just by improving accuracy. Nonetheless, we can, and should, rely on the power of existing predictive methodologies. Through a revised systemic perspective and with appropriate adaptation, conventional learning frameworks should be retooled with a capacity to express and optimize social welfare. This requires providing explicit support for key economics notions: resource scarcity and allocation, human agency and interactions, decision-making, and the role of policy.
We believe it is high time to both confront the challenges and seize the opportunities that learning in social context presents in a disciplined and engaging manner. Welfare considerations now apply to the machine learning community at large, and cannot remain an exclusive niche to work on. Already we are seeing ideas from economics leaking into current machine learning practice (e.g., LLM alignment by preference elicitation); rather than wait for these to surface as grassroots phenomena lacking guidance or a unified conceptualization, we should remain one step ahead and preemptively lay out the necessary groundwork in anticipation of future challenges.
We envision a welfare-centric learning paradigm that is accessible to all, simple to communicate, and appealing to work on—as are most successful learning frameworks. Welfare should be promoted by exploiting the power of learning to obtain accurate predictions, not shying away from it. This can be accomplished by merging the foundational concepts of welfare economics with the established theory and practical know-how of modern machine learning.
Paper outline.
Our main thesis is that machine learning can benefit from the ideas and tools that welfare economics has to offer; Sec. 2 discusses its main principles, and Sec. 3 highlights the potential for synergy with machine learning. In Sec. 4 we present our proposed framework for welfare-maximizing machine learning and details its three orders. Sec. 6 concludes with a look to the future.
2 Welfare economics: crash course
Welfare economics is a field concerned with the characterization, evaluation, and maximization of social welfare in economies and societies. The main principles of welfare economics date back to Adam Smith Adam Smith (1759, 1776), but its formal foundations were laid out only a century later by notable neoclassic economists such as Edgeworth (Edgeworth, 1881), Marshall (Marshall, 1890), Pareto (Pareto, 1906), and Pigou (Pigou, 1920). Modern welfare economists include influential figures such as Hicks (Hicks, 1939), Arrow (Arrow, 1951), Sen Sen (1970), and Stiglitz (Stiglitz, 2012)—all of which received the Nobel Prize for their contributions to this field. Given its rich history, we believe there is much to gain from adopting ideas and perspectives from this well-established discipline.
The distribution of wealth.
The main question welfare economics asks is: how should wealth be distributed across individuals in the economy? Under the working hypothesis that resources (and therefore wealth) are limited, the main object of interest in welfare economics is the Pareto front—the set of all possible economic states in which no individual can be made better off without making things worse for another (see, e.g., Johansson, 1991). In terms of welfare, there are two main considerations for policymakers:
  1. 1.
    Efficiency: How do we reach a Pareto state?
  2. 2.
    Equity: Of all Pareto states, which are preferable?
Efficiency requires an ability to optimize economic outcomes to a point where social benefit is maximal (in the Pareto sense). Markets are a classic example of how the actions of many self-interested agents can combine to produce efficient outcomes (Arrow and Debreu, 1954). However, there are typically many states that are maximally beneficial, but that differ in how benefits are distributed across individuals; in regard to this, markets are mostly silent. As such, equity makes a statement about the relative preference ordering over all possible states, and considers means for steering towards preferable ones. A canonical example is income distribution: all governments likely seek higher overall income (efficiency), but may disagree about whether high income inequality should be permitted or suppressed (equity).
Social welfare functions.
In welfare economics, the primary tool for defining and promoting equity is the social welfare function, which ranks or scores all possible economic states. Welfare functions come in various shapes and colors; for concreteness, we focus on the common choice of cardinal (i.e., real-valued) welfare functions that take the form of an expectation over weighted individual utilities:
${\mathtt{welfare}}({\pi};w)=\mathbbm{E}_{(x,y)\sim{D}}{\left[{w(x)u(x,y;{\pi})}\right]}$ (1)
where ${\pi}$ is a policy (e.g., rules deciding which resources are allocated to whom) and $u(x,y;{\pi})$ is the utility of user $x$ under policy $\pi$, given $y$. Notably, $w(x)$ is a weight function that the planner chooses: this defines the desired direction for overall improvement (i.e., efficiency) by balancing the importances of individual outcomes (i.e., equity). [^2] More generally, weights $w(x)$ can even depend on the utility $u$ of $x$. This allows to express welfare using broader functional forms (e.g., $\min$ instead of sum) or relative measures of inequality; see Appendix A. When the policy ${\pi}={\pi}_{h}$ is guided by a predictive model $h$, we will write ${\mathtt{welfare}}(h)$ to mean ${\mathtt{welfare}}({\pi}_{h})$. See Appendix A for further examples.
The social planner.
Equity is inherently a subjective notion that requires making value judgements. Social welfare functions make it possible to formally express these by setting appropriate weights. In welfare economics, weights are designated by a social planner—either a real or fictitious entity that represents societal preferences. A social planner can set weights to aid low income individuals; implement affirmative action towards some social group; or ensure that all individuals obtain some minimal level of utility. Concrete examples include using weights to prioritize certain subgroups (Björkegren et al., 2022) or balance different objectives (Rolf et al., 2020). The simplest weighing scheme is of course using uniform weights, i.e., $w(x)=1$ for any $x$. But note even this makes a statement, which is that individuals should be weighted by their utility; this is known as ‘utilitarian welfare’. From the perspective of welfare economics, any objective over individuals makes some statement about how value should be distributed.
Human agency.
Welfare economics makes explicit the idea that individuals have agency. Intuitively, this states that individuals (i) want things, (ii) know things, and (iii) act – to get what they want, using what they know. These notions are formally accounted for by modeling utility functions (want), private information (know), and decision-making, either rational or behavioral (act). Any policy that aims to advance welfare must take these into account. Often this requires the planner to make additional efforts, such as to elicit preferences, create incentives for truthful reporting, or infer how users will respond to different policy choices. These are challenging, but give the planner power: if incentives can be aligned, then it becomes possible to harness the willingness of users to invest effort individually for improving outcomes for all; consider public goods and services, crowdfunding platforms, open-source software, and collaborative knowledge bases.
3 Towards welfare-maximizing machine learning
We believe that machine learning has much to gain from adopting a welfare perspective: when inputs represent humans, it becomes possible to promote overall social benefit (efficiency), and imperative to consider how benefits are distributed across the population (equity). We next exemplify this idea on the most basic task of supervised classification, interpreted from a welfare perspective.
Accuracy as social welfare.
Consider a supervised learning task with data sampled as $(x,y)\sim{D}$, and in which inputs $x$ describe individuals (e.g., loan application, resume, movie viewing history, past consumer choices), labels $y$ as describing outcomes of interest (e.g., loan return or default, job qualification, movie rating, satisfaction), and ${D}$ represents the population. The goal is to find a function $h$ from a chosen class $H$ whose predictions $\hat{y}=h(x)$ are accurate on future examples in expectation:
$\operatorname*{argmax}\nolimits_{h\in H}\mathbbm{E}_{{D}}{\left[{\mathds{1}{\{{y=h(x)}\}}}\right]}$ (2)
Because pairs $(x,y)$ are associated with individuals in a population, accuracy admits a welfare interpretation. In particular, it prescribes a specific social welfare function—a special case of Eq. (1)—in which: (i) utility to users derives from accurate predictions, (ii) all users share the same utility function, $u(x,y;{\pi}_{h})=\mathds{1}{\{{y=h(x)}\}}$, (iii) possible outcomes include correct and incorrect predictions, hence (iv) the policy is degenerate, $u(x,y;{\pi}_{h})=u(x,y;h)$, and finally (v) weights are uniform, $w(x)\equiv 1$.
Despite an apparent ‘neutrality’, the above reveals that accuracy maximization in fact makes a particular implicit statement about social preferences. This entails a specific notion of equity—one that derives operationally from the predictive task at hand, rather than from a planned or designated social agenda. But learning is of course not bound to maximizing objectives only of the form of expected accuracy. If our true goal is to maximize welfare, then we should modify the objective to support this goal. For example, if users differ in how much they benefit from accurate predictions, then we can plug in an appropriate utility function $u$; or if there are exogenous constraints on predictions, then we can express these via the policy ${\pi}_{h}$. We will dive deeper into these possibilities in Sec. 4.
3.1 Why welfare?
The above exercise conveys a clear message: once we learn, we are inevitably making some statement about welfare outcomes. This idea applies broadly, including to today’s generative models. A primary benefit of augmenting learning frameworks with a well-defined welfare notion is that such statements can be made explicit. This would not only be an option, but a requirement—working within a welfare-aware framework necessitates making precise statements about efficiency and equity, and makes clear who it is making these statements (e.g., the learner, platform, or government). It also requires to articulate the core components: namely resources, allocation, agency, and policy. Aside from these conceptual benefits, an integrated welfare formalism offers several practical benefits:
Structure.
Welfare economics relies on well-defined social welfare functions, and casts the problem of promoting well-being as one of optimizing such functions. Welfare functions that depend on predictions can naturally be interpreted as learning objectives; in fact, some already are, as we argue in Sec. 4. Given that machine learning has come to excel at optimizing complex objectives over expressive function classes, welfare objectives provide a new target for applying these practices.
Flexibility.
Although we focus here on a particular form, welfare economics is quite flexible in the welfare functions it allows, the types of Pareto frontiers it admits, and the economic structures it supports (Adler, 2019). The formalism is also modular in its main building blocks, such as agents, resources, allocations, decisions, social planner, etc. This aligns well with the modular nature of learning objective in general (e.g., model architecture, loss function, regularization, learning algorithm). It also makes it easy to incorporate some aspects (and not others) when targeting a particular learning task.
Specification.
Accuracy as an objective is notoriously underspecified: for most learning tasks and model classes, there is typically a large set of (approximately) optimal models (Breiman, 2001; Marx et al., 2020; Hsu and Calmon, 2022). These all attain similarly high accuracy but can vary significantly in other important aspects, such as robustness (D’Amour et al., 2022), explainability (see Rudin, 2019), and fairness (e.g., Rodolfa et al., 2020; Coston et al., 2021; Black et al., 2022; Long et al., 2023). Welfare can be useful for breaking ties; in practice, this can sometimes be as simple as applying appropriate regularization (Levanon and Rosenfeld, 2021). We conjecture that in many applications it is possible to improve social outcomes with minimal drop in performance.
Interpretation.
Any learning objective encodes some implicit statement about social preferences; a welfare framework enables to uncover them. It also helps identify components missing in the objective (e.g., agency, interactions), highlight underspecified aspects, and critically examine possible outcomes.
Communication.
When speaking with practitioners, we have found it rather easy to convince that accounting for welfare is important, but quite challenging to explain how it can actually be done. Such efforts require an effective language for communicating novel concepts and complex ideas—which welfare economics provides. Communication is also the reason why our framework has supervised prediction at its core: it is fundamental to learning, widely familiar, and enables us to build around it increasingly complex economic constructs and at different levels of abstraction—namely our ‘orders’.
Impact.
Our final reason for focusing on welfare economics is its proven track record in connecting economic theory, practical econometrics, and real-world policymaking. Machine learning is renowned for its tight and fruitful relations between theory and practice—but has yet to establish an effective interface with policymakers and planners. A welfare perspective can help advance in this direction.
3.2 Alternative views (or: Why not welfare?)
The claim that machine learning should promote social welfare is unlikely to be contested; the debate lies in how to achieve this. Several subfields have made advancements towards particular aspects of this general objective, while others operate on the assumption that accuracy maximization suffices. We challenge these views, and assert that the field at large is still missing a holistic, unifying framework offering actionable guidelines. Here we highlight some alternatives and their connection to our position.
Give accuracy time.
One perspective is that if we give data enough time to accumulate and new methods enough time to improve, then predictive machine learning will organically overcome the concerns we have identified. One example to draw on is how despite many advances in optimization, the simple gradient descent algorithm still drives most modern tools. Another is how large language models have demonstrated that simply predicting the next word with sufficient accuracy and on enough data gives rise to emergent phenomena far beyond this basic task. Perhaps things will just work out?
We are not so optimistic. Our go-to example is the field of recommender systems: Despite a clear incentive to optimize welfare, recommendation systems have for decades settled on accurate prediction of item relevance as a proxy. At the time, and given the inherent difficulties that welfare presents, adopting the probability ranking principle (Robertson, 1977) as a guideline was a reasonable compromise. But over time, we are witnessing the negative implications of prediction-driven recommendations—on individual users, on suppliers and content creators, and on society. Surprisingly (or not), the same philosophy has now come to govern how LLMs are being fine-tuned; consider that methods like RLHF (Ziegler et al., 2019) and DPO (Rafailov et al., 2023) are similarly grounded on inferring user preferences. Predicting "what people want" addresses one important aspect of welfare, but falls short on others. As in recommendation, we conjecture that LLMs will give rise to various forms of scarcity, with competition or conflicts ensuing.
Divide and conquer.
Even if machine learning as a standalone solution does not suffice, one could argue that an economic approach can be applied on top of existing learning tools, rather than integrated within them as we propose. Hence, learning and policy can be advanced independently and combined only later. This is reasonable, and independent efforts and application will likely be required regardless of whether direct integration works or not. But there is increasing evidence that this will not suffice; in fact, the field of fair machine learning rose in response to the clear need for embedding social considerations within the learning objective itself. Advances in the study of fairness in learning have also shown that fairness constraints alone cannot guarantee equity, such as when the effects of learning accumulate over time (e.g., Liu et al., 2018) or as a result of user self-selection (Horowitz et al., 2024). We take these to suggest that interfacing learning and economics requires a holistic approach that enables to anticipate such issues, rather than addressing them separately and in hindsight.
Fairness.
Perhaps closest to our vision, the literature on fairness in machine learning (Dwork et al., 2012) explicitly aims to advance a certain form of social good, and for this it has received much attention. One concern that has been voiced is that fairness focuses on making things equal, but is unable to express the benefit that results from such efforts (Heidari et al., 2018; Hu and Chen, 2020; Kasy and Abebe, 2021). Some initial works have proposed welfare as an alternative (Shirali et al., 2024; Perdomo, 2024), but have yet to gain much influence. From our perspective, fairness takes one step towards acknowledging that something is of limited supply (typically positive predictions), but this is rarely made explicit. More importantly, fairness constraints do not account for agency: they do not ask what users want, nor how much they stand to lose or gain. Fairness constraints also do not anticipate how users will act; for example, predictions for school admissions may satisfy demographic parity, but if members of the minority group do not choose to apply initially, then this guarantee is broken. Other than some notable exceptions (e.g., Kannan et al., 2017), works on fairness generally do not consider (un)truthful reporting, manipulation, or information asymmetry—all central aspects of economic agent modeling. Whereas fairness typically aims to "solve" inequity, welfare provides a framework for advancing flexible goals.
Strategic learning.
Another relevant line of research is that of strategic learning, such as strategic classification (Brückner et al., 2012; Hardt et al., 2016) and performative prediction (Perdomo et al., 2020), which has drawn much recent interest. Strategic classification considers agency, but mostly in the form of “gaming”, and does not model interactions despite its canonical examples (e.g., admissions) having inherent constraints (e.g., quotas) (Liu et al., 2022). Some works discuss social aspects of learning (e.g., Milli et al., 2019), but the primary focus remains to be on maximizing accuracy. Performative prediction defines optimality and stability only w.r.t. the loss function, and so performatively optimal solutions can in principle have arbitrarily low welfare. Welfare is not a driving consideration, nor is it an integral part of current approaches. Focus in both fields has so far been predominantly theoretical, with little work done on practical tools and empirical analysis.
Synergy.
Welfare presents some inherent challenges for learning that may initially seem discouraging. We propose to draw inspiration from other fields that have successfully contended with similar or comparable challenges. Adversarial learning (Goodfellow et al., 2014) is a prime example for a field that has transformed an inherently game-theoretic problem to an engaging and highly relevant applied task. It’s worst-case perspective, however, can be very limiting, and the field may benefit from finer modeling of incentives and resources (Ehrenberg et al., 2025). Decision-focused learning (Mandi et al., 2024) and algorithms with predictions (Mitzenmacher and Vassilvitskii, 2022) have successfully built on predictive machinery to solve complex downstream tasks, as has the emerging field of prediction-powered inference (Angelopoulos et al., 2023) for improving statistical inference. A welfare framework can provide these fields with new economics-driven challenges. Both agency and resource limitations can often be cast as particular distribution shifts; distributional robustness (Chen et al., 2020a) and out-of-distribution generalization (Bühlmann, 2020) can provide tools, as well as benefit from new technical challenges.
Welfare without welfare.
Welfare economics is not the only approach for reasoning about and facilitating welfare, nor is it free from issues and limitations in itself. Criticism includes its subjective nature; the need to measure and compare utility across individuals; the emphasis on cardinal rather than ordinal utilities; the reliance on assumptions of rational behavior; the susceptibility to externalities and other sources of market failure; the need for a centralized social planner entity; and challenges in policy evaluation. Other schools of thought in economics offer alternatives: For example, Sen’s capabilities approach (Sen, 1999) focuses on ensuring people are capable of achieving what they seek, rather than the value of what they obtain. Within machine learning, there been have calls for alternative approaches as well, such as to ‘democratize’ the issue of alignment using social choice theory (Conitzer et al., 2024; Ge et al., 2024; Fish et al., 2024). We view these as complementary to ours, and believe there is merit in advancing welfare in machine learning simultaneously along several fronts.
3.3 From accuracy to welfare: paving the path
In courses and textbooks, machine learning is classically portrayed as a closed system: inputs go into a machine which then generates outputs. But even simple correlation machines that are deployed in the real world become part of it. For social settings comprising many other agents—namely us—we believe there is need for a refined systemic perspective. As the above examples convey, steps have been taken in various directions; these are praiseworthy, but lack unifying principles and guidance.
Although our goal is to discuss how learning can generally promote welfare, we believe there is value in keeping accuracy an integral part of the learning objective. This is both because accuracy has value in and of itself, and because predictive methods are, and will likely remain, a primary tool used in practice. We therefore argue for the deliberate extension of conventional supervised learning, not its replacement. Supervising learning already proposes various tools that can be harnessed, such as imposing cardinality constraints (Lapin et al., 2015; Petersen et al., 2022), controlling error types (Menon et al., 2019), or cost-sensitive optimization (Elkan, 2001). It also offers effective means for advancing general research such as benchmarks and competitions.
In Sec. 4 we present a proposed framework for welfare-aware learning that gradually extends from basic accuracy maximization towards increasingly complex economic objectives. One benefit of such a bottom-up framework is that it enables researchers and practitioners to choose the level of abstraction they wish to work in, such as assumptions on behavior, range of possible incentives, dependencies across agents, the role (or existence) of a central planner, and the type of pursued welfare objective. We believe this can make welfare-aware machine learning an engaging and appealing topic to work on for people of diverse backgrounds, interests, knowledge, and orientation.
Who drives welfare?
In the late 1940s, the field of economics found itself gaining significant social power, with its newly developed theoretical ideas and practical tools having unprecedented impact on society. Eighty years later, it appears we are witnessing similar trends for machine learning; it therefore only makes sense to try and learn from these earlier attempts—both successful and less so. One conclusion is that the wheel need not be (re)invented, only adapted. This is one of the grounds for our argument of drawing on ideas from welfare economics. But a major lesson learned is that economic objectives, welfare included, can rarely be forced upon a system. Economics views welfare as an emergent outcome arising from the decentralized interactions of many agents, particularly in idealized settings such as markets with perfect competition; this is the ‘invisible hand’ in essence. But in realistic setting, welfare needs guidance to materialize, whether through legislation, incentive alignment, or even minor nudges. As researchers, our goal should be to understand how systems can be steered towards beneficial outcomes, and what market failures can occur absent such direction. As system designers, we should carefully avoid coercing our own beliefs, and instead enable policymakers to implement their policies through the tools we build while maintaining transparency and accountability.
Who should care?
We believe the following different populations of researchers, with different levels of familiarity to welfare in machine learning, may all find our narratives helpful. The first, and perhaps foremost, is the general ML researchers and practitioners working on problems that have downstream social impact, [^3] Clear examples are people working on recommendation, social networks, and e-commerce. But we anticipate that over time, more and more subfields of ML will find themselves in a position where their work begins to have concrete social impact; consider text and image generation as recent examples. It is crucial to think one step ahead before we start to face serious welfare issues. but who do not account for this explicitly. Here our goal is not only to raise awareness to welfare issues, but also to provide a tangible means towards achieving it. This is why our framework below has supervised prediction at its core: it is fundamental to ML, widely familiar, and enables us to build around it increasingly complex economic constructs and at different levels of abstraction–namely our ‘orders’. The second population features researchers in related sub-fields such as fairness (Dwork et al., 2012) or strategic classification (Hardt et al., 2016), which is a considerably smaller group within the entire ML community. Here we intend to provide a broader framework as well as complementary perspective to the questions they are already pursuing. For example, fairness research has mostly overlooked agency, whereas strategic classification rarely considers resources. By revealing such “blind spots” we wish to motivate researchers to confront them. Finally, the third population consists of researchers already working at the intersection of ML and (welfare) economics. This is perhaps an even smaller group within the general ML community (for which utility and welfare have been overlooked). Our goal here is to provide an organizing conceptual framework that is firmly grounded in the core concepts of machine learning.
4 Proposed framework
Figure 1: Proposed framework: The three orders of welfare-maximizing machine learning.
To concretize our position, we propose a formal framework for welfare-maximizing learning, organized into three hierarchical ‘orders’. Our framework presents a hierarchy of problem formulations for welfare-maximizing learning, organized into three ‘orders’ of gradually increasing complexity. The orders are built in a bottom-up fashion: The lowest order (Order 0) coincides with the standard objective of maximizing predictive accuracy, but provides a novel welfare perspective suited for social tasks. Each higher order then builds on and generalizes the one below it, this by adding to the objective another layer of economic complexity: first focusing on system decisions (Order 1), and then on user choices (Order 2). As such, our framework is based on accuracy maximization at its core, but enables—and in fact requires—to explicitly model resources, allocations, and agency. It also requires to specify the role predictions play in shaping social outcomes. An illustration of the framework and its orders is given in Fig. 1.
We view our framework as a means to enable the effective communication of ideas regarding welfare to different target audiences. For example, we have found it rather easy to convince ML practitioners that welfare is important, but quite challenging to explain how this can be done in practice. We envision our framework as enabling the flow of ideas and methodologies between ML and welfare economics in both directions: Economics offers many compelling ideas, but interfacing them with ML is rarely straightforward. Conversely, ML offers many powerful tools, but these have yet to be harnessed for maximizing welfare.
4.1 Order 0: Allocating Accuracy
Consider first the standard accuracy objective in Eq. (2), and its interpretation as a welfare objective in which users gain utility from obtaining accurate predictions. Accuracy will be beneficial to users in any a platform or service that relies on predictions; consider recommendation systems, online market platforms, financial aid tools, or diagnistic health services. We argue that in such cases, and as long as learning is prone to some level of error (which is plausible), then accuracy itself is a limited resource—simply because not all users can obtain the same level of accuracy. Thus, the choice of classifier becomes, in effect, a decision on which users will be ‘allocated’ more accurate predictions, and which less. This then begs the question: who should be allocated an accurate prediction? One answer is users who most need it, or who benefit mostly. In the reasonable case where there is natural variation in individual utilities, we can to optimize a utility-weighted accuracy objective:
$\operatorname*{argmax}\nolimits_{h\in H}\mathbbm{E}_{{D}}{\left[{u(x,y;\hat{y})\cdotp\mathds{1}{\{{y=h(x)}\}}}\right]}$ (3)
where $u(x,y;\hat{y})$ is the utility for user $x$ given prediction $\hat{y}$. This transforms Eq. (2) into an explicit utilitarian welfare objective in which individualized utilities derive from prediction correctness: $u(x;{\hat{y}})$ if ${\hat{y}}=y$, and zero if not. Conditioning $u$ also on predictions ${\hat{y}}$ enables to express utilities that depend on predictions themselves (beyond their correctness); for example, if users benefit more from positive predictions, as in loans, hiring, admissions, etc.
Agency.
While apparently simple, optimizing Eq. (3) poses significant challenges once we accept that users have agency. If users seek to promote their own interests, and have control of information that is important to the system, then their choices regarding how and what to report can significantly affect learning outcomes. For example, consider that if we simply ask people to report their utilities, then each individual will be incentivized to report the highest possible value, making all reports completely uninformative. Another example is that users may be prone to manipulate their features if this helps them secure favorable predictions, as is the case of strategic classification (Hardt et al., 2016).
Challenge: Optimizing welfare objectives.
Maximizing welfare through learning presents new algorithmic challenges. One task is to identify learners that optimize a given social welfare function—intrinsically, a question of identifying a desirable point on the Pareto front. Another task is to generate all points on the frontier (see, e.g., Navon et al. (2021)). One possible approach is to cast Eq. (3) as a problem of distribution shift. Here shift can stem from several factors, including the welfare function weights, user utilities, and user inputs. But agency means that utilities can be misreported and features manipulated; as a result, the way in which the distribution shifts becomes dependent on the choice of classifier, indirectly through how users respond. This makes Eq. (3) a challenging instance of model-dependent distribution shift (Drusvyatskiy and Xiao, 2023). Here the interaction between the system and its users can be modeled as a Stackelberg game (Chen et al., 2020b). This perspective has been useful for questions on learnability, e.g. via generalization bounds that rely on strategic VC analysis (Zhang and Conitzer, 2021; Sundaram et al., 2023) or that accommodate feature-dependent utilities (Pardeshi et al., 2024). Nonetheless, many important open questions remain.
4.2 Order 1: Incorporating System Decisions
Generally, predictions are useful for the system if they aid in making better decisions about uncertain outcomes. Since decisions in social settings often directly prescribe how to allocate limited resources to users, a useful next step is to encode them explicitly into the objective. We formalize this idea by modeling a system that makes decisions about users (e.g., who to hire) through predictions (e.g., resume screening). Consider a user $x$ with label $y$, and denote by $a$ the action the system takes (e.g., hire or not). Let ${r}(a,y)$ be the reward for the system on this user given action $a$ (e.g., the quality of the candidate, if hired). The direct dependence of ${r}$ on $y$ means that if we know $y$, then we can write the optimal action as $a^{*}={\pi}(y)$ for some policy ${\pi}$. Since $y$ is generally unknown, the common approach is to replace $y$ in ${\pi}$ with a prediction ${\hat{y}}=h(x)$, denoted $a={\pi}(h(x))={\pi}_{h}(x)$, in hopes that better predictions translate to better decisions. Such policies, referred to as prediction policies (Kleinberg et al., 2015), are appropriate when the uncertainty in $y$ is a stronger factor for outcomes than potential causal effects (we discuss this distinction further in Section 5.4). In this setting, choosing to work with prediction policies gives reason for the system to optimize $h$ for accuracy (Eq. (2)).
From predictions to actions.
We consider cases where there is a global restriction on the set of all actions $a$. These can express, for example, a limited number of available jobs (via cardinality constraints), a total sum of funds that a bank can lend (knapsack constraints), regulation on the amount of financial risk an insurer can take (bounded expected risk), or a limit on the number of posts a social platform can block to still enable free speech (lower-bounded rates). Given a set $A$ of feasible actions, the objective can be written as:
$\operatorname*{argmax}\nolimits_{h\in H}\mathbbm{E}_{{D}}{\left[{{r}({\pi}_{h}(x),y)}\right]}\quad\text{s.t.}\quad{\pi}_{h}({D})\in A$ (4)
where ${\pi}_{h}({D})$ is the set of actions over the entire population.
Agency.
Machine learning has many tools for coping with constraints. For example, if there is a quota on the number of available interview slots, then top-$k$ classification (e.g. Lapin et al., 2015; Petersen et al., 2022) or ranking (e.g. Cao et al., 2007) seem like adequate solutions. The crux, however, is that such methods do not account for agency: since users seek to be ‘allocated’ (e.g., get the interviews), they will likely try to present themselves as relevant (e.g. by tweaking their resume). A key point is that since users contend over resources, this introduces dependencies into their actions and subsequent outcomes. Such behaviors are ubiquitous in economics (e.g., costly signaling, adverse selection), but have been underexplored in machine learning.
Social welfare.
In terms of welfare outcomes, we make a distinction between two settings of interest:
  • An aligned setting in which ${r}={\mathtt{welfare}}$, and so the interests of the system align with societal preferences, e.g., as for government and nonprofit organizations. Note Eq. (3) becomes a special case of Eq. (4) when utility derives from accuracy and $A$ does not impose restrictions.
  • A misaligned setting in which ${r}$ can be at odds with welfare, for example if it concerns revenue or user engagement, as is more common in commercial settings. Here a social planner is needed to incentivize the learner to account for welfare outcomes, as we detail in Sec. 5.1.
Challenge #1: Decisions and externalities.
At Order 0, user responses are generally made independently, and so the learning objective decomposes over examples. [^4] Note this is a key assumption in standard strategic classification. But once the system makes decisions under constrained resources, user behavior can become intricately dependent. Consider for example admissions under a limited quota: here applicants cannot simply pass a fixed acceptance threshold, but instead, must surpass other applicants—i.e., the bar adjusts according to competition. The tension for both decision-makers and applicants tends to focus on ‘marginal’ students; complex gaming behavior among such candidates is well-documented in the education literature (Bound et al., 2009). In economics, inter-user dependencies are known as externalities. One way to model how learning creates externalities and coordinates user behavior is by casting the problem as a Stackelbeg-Nash game (Liu, 1998), where the system plays first, and users respond collectively by playing an induced simultaneous game (with externalities). Not much work has studied learning in this challenging setting.
Challenge #2: Misaligned system and user objectives.
As we note, generally we cannot expect the reward for the system to align with user interests. If no external forces intervene, then the challenge is to learn a decision policy that is effective at the Stackelberg-Nash equilibrium. This necessitates learning objectives that can express and anticipate how users will collectively respond to a deployed model. Contrarily, if there is regulation (e.g., on welfare or risk), then the system’s decision space becomes much more limited. An interesting question then is how to align incentives through regulation. For example, would requiring the system to optimize a weighted sum of accuracy and welfare help to promote equity? And if so, which weights, and by what means?
4.3 Order 2: Enabling User Choices
Even if a system has the capacity to make decisions about users, often those very users will also have some say regarding final outcomes. For example, even in hiring and admissions, candidates must first choose to apply, and if selected, choose to accept. Thus, both reward for the system and utility to users depend on outcomes that result from the interplay of system decisions and user choices. Similarly to system decisions, we will model user choices as also depending on predictions, or more generally on the learned model $h$. Let ${c}_{{\pi}_{h}}(x)$ be the choice of user $x$ in response to $h$. Incorporating user choices into the system’s objective gives:
$\operatorname*{argmax}_{h\in H}\mathbbm{E}_{{D}}{\left[{{r}({\pi}_{h}(x),{c}_{{\pi}_{h}}(x),y)}\right]}\quad\text{s.t.}\quad{\pi}_{h}({D})\in A$ (5)
where ${r}$ now depends jointly on system decisions ${\pi}_{h}$ and user choices ${c}_{{\pi}_{h}}$ through the learned predictor $h$.
Agency and resources.
When users make choices, this makes them a part of the allocation process—but also the reason for scarcity. Here we outline three scenarios of interest:
  • Self-selection: For most systems that makes decisions about users, those users must first choose to join the system; consider hiring, admissions, medical programs, educational aid, and welfare benefits. In econometrics, the choice of users to participate (or not) based on the utility they expect to obtain is known as self-selection (Roy, 1951). Since self-selection in aggregate determines the user population, once user decisions depend on the learned model, learning obtains the power to shape the population’s composition—a power which should not be taken lightly. Although self-selection has a long history in economics, it is only now beginning to draw attention in machine learning research (Zhang et al., 2021; Cherapanamjeri et al., 2023; Horowitz et al., 2024).
  • Matching: Prime examples in which user choices drive economic outcomes are recommendation systems and online market platforms. These typically include a single learner whose role is to match users with relevant items (Jagadeesan et al., 2023a); once matched, it is the choices of users that ultimately determine allocations. Since matchings are often based on predictions, their quality determines outcomes for users, suppliers, and the system. This makes them unique in relation to conventional two-sided markets (Rochet and Tirole, 2003). In modern human-facing systems, key resources primarily include the attention, time, and budgets of users. This calls for careful and responsible modeling of user choice behavior (Kleinberg et al., 2024).
  • Competition: For many tasks and services of interest, users can now choose between multiple platforms or providers. This creates competition: once providers aim to maximize their market share, users themselves become the scarce resource, over which providers compete (Ben-Porat and Tennenholtz, 2019; Jagadeesan et al., 2023b; Yao et al., 2023, 2024). When the benefit of users depends on the quality of predictions, competition will revolve around which provider learns more accurate models—and for which users. This creates a novel ‘market for accuracy’, in which learning can have much impact on efficiency and equity.
Social welfare.
A canonical property of classic markets is that they coordinate the behavior of many self-interested agents in a way which can naturaly lead to welfare maximization (e.g., Arrow and Debreu, 1954; Shapley and Shubik, 1971). A key question in our context is whether this emerges also in markets where coordination is mediated by predictions (Nahum et al., 2024). Given the growing concerns regarding how recommendation systems may drive polarization, echo chambers, informational barriers, economic inequity, and unhealthy usage patterns—the answer is likely negative. Welfare economics may be helpful in posing the question of why this happens as one of market failure. This provides tools for uncovering the mechanisms underlying failure, such as negative externalities, public goods, information asymmetry, market power and control (e.g., by monopolies), or collusion, as they manifest through learning.
Challenge #1: Escaping echo chambers.
Since its early days, the task of recommending content has been treated as a pure prediction problem. But accumulating evidence of its likely ill effects has motivated a search for more viable alternatives. A major issue is the reinforcing nature of accuracy-driven recommendation: users that are recommended certain items become ‘locked in’ on similar content through a positive feedback loop of indefinite model retraining, choice behavior, and recommendation. One aspect of this loop that is often overlooked, and relevant to our context, is that recommendations also incentivize the creation of new content by exposure-seeking creators. Recommendation is essentially a problem of matching demand (of users) and supply (of new content) over a bi-partite user-content graph; thus, if we wish to prevent the formation of echo chambers and filter bubbles, we should find ways to exploit this structure to promote socially-beneficial outcomes. Some examples include incentivizing exploration (Mansour et al., 2020) or the creation of diverse content (Eilat and Rosenfeld, 2023; Yao et al., 2024), but many other paths are possible.
Challenge #2: Learning to compete.
Once predictions are given as a service, it is only natural that multiple providers will compete over who can give users better predictions. One example are housing market platforms such as Zillow and Redfin which compete over who provides better pricing recommendations. To some extent, another example are LLMs who compete over better text generation. Competition is driven by the fact that users are free to choose a provider, but that choices are costly (e.g., time, premium features, opportunity costs). Thus, although users gain from accuracy, optimizing expected accuracy is likely not a good strategy for providers. As a result, welfare could suffer. Welfare in such markets will be high if the market is efficient. This requires a planner that can steer competing learners towards a favorable Nash-Stackelberg equilibrium, where users follow the simultaneous moves of multilpe systems.
5 Further Prospects and Challenges
Sec. 4 lays out the main goals and principles of our framework, along with possible avenues to pursue within each of its three orders. But there are further opportunities, as well as challenges, that lie beyond the framework’s core and are of both interest and significance. In this section we highlight some of these directions.
5.1 Regulation: Social planner, revisited
Our perspective of a social planner so far has been that of a useful construct for defining the social preferences that should guide learning. As such, we have mostly discussed the question of how to learn in a way that promotes or aligns with given social preferences. In a sense, this depicts the social planner as acting before learning. But there are additional roles a social planner can play during or after learning, actively and through various mechanisms. For example:
  • Monitoring. One defining characteristic of data-driven systems is that they hold a distinct informational advantage over their users. In economics, information asymmetry is a well-known source of hazard for markets, with two prominent issues known as moral hazard (Holmström, 1979) (arising from asymmetric information on decisions) and adverse selection (Akerlof, 1978) (arising from asymmetric information on system states). Monitoring can help to mitigate this by identifying what information gives the system an unfair advantage, and then channeling this information to users. This provides an operational motivation for transparency, with the definitive purpose of improving welfare.
  • Subsidizing. When resources are limited, their allocation often depends on the wealth or assets of individuals. This presents an opportunity to intervene by injecting funds into the system in a way that balances outcomes, e.g., via targeted subsidies. Subsidies can either be given in response to a learned classifier, or determined by policy prior to learning. Since outcomes in our case depend on predictions, this presents a unique opportunity for a social planner to intervene by directly providing users with useful ‘features’, thus bypassing the need for a monetary mechanism. An interesting question is then how to design learning objectives that operate under such subsidies.
Other roles include: setting standards, licensing, auditing, mandating disclosure and transparency, and stress-testing.
5.2 Inverse tasks: policy preference elicitation
A key point we have made throughout is that any prediction policy advances some social preferences. Until now, our focus was on learning to maximize welfare under a specified social welfare function. But we can also ask the inverse question: given a policy—what social welfare function is it optimizing? This can be useful for revealing the preferences of general, unspecified policies, and for evaluating whether a policy that has been specified truly achieves its goals. One way to approach this would be to use tools from econometrics, e.g. as in (Björkegren et al., 2022; Pardeshi et al., 2024). But is also interesting to ask whether ideas from inverse problems in machine learning can be adopted.
5.3 Evaluation
A main driving force of modern machine learning is the ability to evaluate the performance of different methods effectively and consistently under a wide range of settings and conditions. Unfortunately, such privileges are not possible when learning in social settings. The challenge of evaluation is shared by all fields at the intersection of learning, economics, and human behavior, but welfare-maximizing machine learning presents its own unique difficulties. For example, even the basic task of measuring or estimating social welfare is far from trivial. One immediate challenges is the difficulty (and restrictions) of obtaining and working with representative and informative user data. Another is that, as we have stressed, welfare depends on outcomes, which are the product of decisions due to policy, and so evaluation requires either control of the policy or a way to estimate counterfactuals. Other than a few noteworthy exceptions (e.g., Björkegren et al., 2020; Haupt et al., 2023; Mendler-Dünner et al., 2024), most current works settle for ‘semi-synthetic’ evaluation that uses real data and simulated user behavior. Simulation is often useful as an intermediary step to drive the field forward (e.g., as in reinforcement learning), but the ultimate goal should be to enable seamless, realistic, in-the-wild evaluation.
5.4 Accounting for causal effects.
The framework we presented makes use prediction policies (Kleinberg et al., 2015) that specify the dependence of actions on predictions as $a={\pi}_{h}(x)={\pi}(h(x))$. The key assumption of prediction policies is that knowing the true $y$ suffices for making an optimal decision; hence the motivation for replacing $y$ with the prediction ${\hat{y}}=h(x)$. But in realistic settings, $y$ is rarely the only source of uncertainty. Most decision settings have a casual aspect, meaning that actions $a$ can affect outcomes $y$. Hence, there is no ‘true’ $y$ on which we can draw, and the optimal action $a$ (or ‘intervention’) must include an understanding of its causal effects on $y$.
When learning in social settings, the issue of causality is further complicated by the fact that users have agency, and therefore act. Because learning determines policy, and policies shape users’ choice behavior, learning comes to have a second-order causal effect on $y$. The task of causal inference therefore attains two goals: understand the impact of decisions on outcomes directly, and indirectly through user choices.
Econometrics offers many tools for contending with questions of causal estimation. The main challenge is to make inference of causal effects on the basis of observed data—which is the common setting considered in learning. In cases where we can experiment, randomized control trials (or A/B tests) offer an alternative. Some of these ideas have been used in the literature on strategic learning (Miller et al., 2020; Shavit et al., 2020; Horowitz and Rosenfeld, 2023), mostly in relation to the interaction between the system and its users. A prominent point that concerns welfare is that a causal understanding can be used to improve outcomes—if incentives are set correctly (Kleinberg and Raghavan, 2020). Despite the importance of causal effects, we believe there is merit in first exploring question of welfare under policy predictions, as means to set the stage for more advanced questions concerning causality.
6 Discussion
Economics studies the tensions that arise from having limited means but unlimited wants. Our main thesis is that while the appetite for what machine learning can give us only grows, we do not yet have the tools to integrate the limitations inherent to any social system in a way that supports informed policy choices. Welfare economics provides general guidelines for designing systems that enable a social planner to express and promote outcomes that align with societal preferences, as they relate to scarcity in resources. We have identified ways in which we believe this perspective applies to, and can be useful for, the design of learning systems that operate in social contexts. Scarcity in learned systems is often far more illusive than in classic economic settings that consider tangible goods. But nonetheless, it exists, and its implications bear much significance for the value that we can hope to obtain from using machine learning as individuals. Our proposal to augment conventional learning frameworks with a notion of social welfare serves as a first step towards this goal. But the road ahead is long and challenging, and requires careful deliberation of the societal role we envision for machine learning.
Acknowledgments
Haifeng Xu is supported by the AI2050 program at Schmidt Sciences (Grant G-24-66104), Army Research Office Award W911NF-23-1-0030, and NSF Award CCF-2303372. Nir Rosenfeld is supported by the Israel Science Foundation grant no. 278/22.
Appendix A Example social welfare functions
Welfare economics considers a wide array of social welfare function types. Here we survey a small subset of the most popular alternatives. The primary distinction is between ordinal functions, which rank all possible economic states in terms of their social preference, and cardinal functions, which assign particular scalar values to states. While any cardinal function induces an ordering, instantiating such functions typically requires access to (and the existence of) individual utility values, which are not always observed (or even attainable) and can be compared on a single shared scale (e.g., monetary value). This is why much of the literature considers cardinal functions, as we do here.
The simplest form of social welfare function considers the average utility over all individuals in a society:
$\frac{1}{n}\sum_{i=1}^{n}u_{i}$
which is known as utilitarian welfare. This is commonly generalized to include (i) a mapping $g$ of individual utilities $u_{i}$ to socially relevant values, and (ii) a weighting scheme $w$ that assigns different weights to different cases. The general form is:
$\frac{1}{n}\sum_{i=1}^{n}w_{i}g(u_{i})$
where $g(u)=u$ and $w_{i}=1$ recovers pure utilitarian welfare. A common choice is setting $g$ to be concave to encodes risk aversion. Weights can more generally be functions (rather than fixed scalar), and both $g$ and $w$ can depend on individual utilities or on the set of all utilities, on individual properties (features) $x_{i}$, and on additional exogenous variables $z$. Thus, despite its simple form, average welfare can be a rather flexible notion. These extensions apply to all social welfare functions described below, and so for simplicity we will present them in the above format.
An alternative to prioritizing is the egalitarian approach which instead aims to maximize welfare for the worst-off individuals in society. This is known as Rawlsian welfare, given by:
$\min\nolimits_{i\in[n]}w_{i}g(u_{i})$
Note this can be expressed also by an average welfare function with utility-dependent weights defined as $w_{i}(u)=\mathds{1}{\{{i=\operatorname*{argmin}_{j}g(u_{j})}\}}$ for $u=(u_{1},\dots,u_{n})$. This suggests a smooth transition between the two types using softmin weights, $w_{i}\propto e^{-\beta g(u_{i})}$, which as a vector approaches the argmin indicator for $\beta\to\infty$. Less common, but possible, is to replace the min with a max to promote the most well-off individual (with no consideration of others); this applies for example to settings in which different societies compete in winner-takes-all matches, and the goal is to advance the best possible candidate (e.g., international sports competitions).
Other common alternatives exist for interpolating between the utilitarian and Rawlsian approaches. One approach is to consider a convex combination of both, namely:
$\alpha\frac{1}{n}\sum_{i=1}^{n}w_{i}g(u_{i})+(1-\alpha)\min_{i\in[n]}w_{i}g(u_{i})$
for some $\alpha\in[0,1]$. A second is to use geometric averaging, known as Nash welfare:
$\left(\prod_{i=1}^{n}w_{i}g(u_{i})\right)^{\frac{1}{n}}$
The third and most flexible approach is the iso-elastic parameterization:
$\frac{1}{n(1-\gamma)}\sum_{i=1}^{n}(w_{i}g(u_{i}))^{1-\gamma}$
which recovers utilitarian welfare for $\gamma=0$, Nash for $\gamma\to 0$, and Rawlsian for $\gamma\to\infty$.
Social welfare functions can also be made to explicitly account for inequality. For example, Sen’s welfare function is $W(u)(1-G(u))$ where $W$ is any base welfare function, and $G$ is the Gini inequality index. Foster’s welfare similarly considers $W(u)e^{-T(u)}$ where $T$ is the Theil index.